OpenStudy (anonymous):

After 1 minute, a file is downloading from the internet is at 0 kilobytes downloaded. After 2 minutes, the file is at 7 kilobytes downloaded. After 3 minutes, the file is at 26 kilobytes downloaded. After 4 minutes, the file is at 63 kilobytes downloaded. Write an expression to model the amount of the file downloaded

OpenStudy (anonymous):

m = number of minutes y = number of kilobytes downloaded y = m^3 - 1 Try it out. 2 minutes = (2^3) - 1 At 2 minutes, 7 kilobytes is downloaded. It works for 3 minutes and 4 minutes too. Reasoning? I just looked at the numbers and noticed it must be some kind of exponential function (it's increasing at a huge rate!)

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

quick question. where did the root3 and -1 come from?